OpenStudy (xapproachesinfinity):

Hi guys, what do you think about this problem if differentiate \(\huge \rm (x^2)'=2x\) if i think of \(\huge \rm x^2\) as \(\huge \rm x.x=(x+x+x+x............x)\) and then I differentiate it \(\large \rm (x+x+x+....+x)'=1+1+1+.....+1=x\) what is the problem here?

OpenStudy (xapproachesinfinity):

it is legit that differentiating the sum is the sum of derivative! so where is the issue

OpenStudy (xapproachesinfinity):

@ganeshie8

OpenStudy (ikram002p):

wow its really something wrong :P

OpenStudy (xapproachesinfinity):

Have you ever thought about it? it is interesting thing

OpenStudy (ikram002p):

i thought about it :O i said something wrong xD

OpenStudy (xapproachesinfinity):

oh! so what did you think was wrong here?

OpenStudy (xapproachesinfinity):

:P lol it a bit complex how things get to this but it must be something not really that big is the issue here

OpenStudy (michele_laino):

derivative function as you have applied, is defined for a real and complex functions only!

OpenStudy (xapproachesinfinity):

it is true x.x=x+x+x+...+x x times

OpenStudy (xapproachesinfinity):

according to our arithmetic laws

OpenStudy (ikram002p):

ok , X is variable how u can be sure it applies the sum of x times ? it might be some negative number so let x=-3 x^2=-3*-3 can u express this in sum xD u have to take |x| lol sort of creepy

OpenStudy (michele_laino):

whereas, in your second line, you treat the function y=x*x as a integer function

OpenStudy (xapproachesinfinity):

this is also real @Michele_Laino so that's not the issue x^2=x.x not problem about that

ganeshie8 (ganeshie8):

this is one of the classic fake proofs for proving 1=2

OpenStudy (ikram002p):

yeah , nice proof though

OpenStudy (xapproachesinfinity):

No x variable can be written as x^2 can be written as x.x or any other form no problem as far as i know

OpenStudy (ikram002p):

ok write this in arithmetic way (1/2)^2

OpenStudy (xapproachesinfinity):

@ganeshie8, my prof give us this! he said it is not the type of proving 1=2

OpenStudy (xapproachesinfinity):

1/2.1/2? the same as 1/4

OpenStudy (michele_laino):

If x^2 is a real function, for example, you can not write x*x=x+x+...x, because x is a real number, and the set of real number is not numerable.

OpenStudy (ikram002p):

ok now lets this of the differential definition hmm

OpenStudy (xapproachesinfinity):

the prof said it is about how we differentiated the x+x+x+...+x

OpenStudy (anonymous):

can you please come and help me on a problem

OpenStudy (ikram002p):

y=(x^2) dy/dx=2x x.x=x+x+x+x...+x ( x times ) think of it like (gf)' xD

OpenStudy (xapproachesinfinity):

So the LHS is 2x but the right hand side is what?

ganeshie8 (ganeshie8):

yeah that will work, then use the product rule

OpenStudy (michele_laino):

@xapproachesinfinity If x^2 is a real function, for example, you can not write x*x=x+x+...x, because x is a real number, and the set of real numbers is not numerable

OpenStudy (xapproachesinfinity):

@Michele_Laino I said there is no problem whatsoever about how we know a function this is a differentiation problem

OpenStudy (xapproachesinfinity):

so the LHD is good! no problem about it the right hand side causes the problem lol something is wrong we we differentiate that sum

OpenStudy (xapproachesinfinity):

Oh great! I guess i kind of see the flaw of this problem